# Online Posterior Sampling with a Diffusion Prior

### v1

@@ 4 Posterior Sampling
This section is organized as follows. We derive a general posterior distribution for a chain of random variables conditioned on observations in Section 4.1. After that, in Sections 4.2 and 4.4, we specialize it to the observation models in Section 2.
@@ 4.1 Chain Model Posterior
Let $h=\left\{(\phi_{\ell},y_{\ell})\right\}_{\ell\in[N]}$ denote a history of $N$ observations (Section 2) and $H$ be the corresponding random variable. In this section, we assume that $h$ is fixed. The Markovian structure of the reverse process in Figure 1 implies that the joint probability distribution conditioned on $h$ factors as
@@
Let $p$ be a probability measure over the reverse process (Figure 1). Then
@@
The claim is proved in Section A.1. ∎
@@ 4.2 Linear Model Posterior
Now we specialize Lemma 1 to the diffusion model prior (Section 3) and linear models (Section 2.1). The prior distribution is the reverse process in (5),
$p(s_{T})=\mathcal{N}(s_{T};\mathbf{0}_{d},I_{d})\,,\quad p(s_{t-1}\mid s_{t})=\mathcal{N}(s_{t-1};\mu_{t}(s_{t}),\Sigma_{t})\,.$
The term $p(h\mid s_{0})$ is the likelihood of observations in Assumption 1. To fully instantiate Lemma 1, we make additional assumptions on the integrals in it.
@@
Let $p$ be a probability measure over the reverse process (Figure 1). Let $\bar{\theta}$ and $\bar{\Sigma}^{-1}$ be defined as in (3). Suppose that we approximate
$\textstyle\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{t})\mathop{}\!\mathrm{d}s_{0}\approx p(h\mid s_{t}/\sqrt{\bar{\alpha}_{t}})$ (6)
for all $t\in[T]$. Then $p(s_{T}\mid h)\approx\mathcal{N}(s_{T};\hat{\mu}_{T+1}(h),\hat{\Sigma}_{T+1}(h))$, where
[…]
For $t\in[T]$, we have $p(s_{t-1}\mid s_{t},h)\approx\mathcal{N}(s_{t-1};\hat{\mu}_{t}(s_{t},h),\hat{\Sigma}_{t}(h))$, where
$\hat{\mu}_{t}(s_{t},h)=\hat{\Sigma}_{t}(h)(\Sigma_{t}^{-1}\mu_{t}(s_{t})+\bar{\Sigma}^{-1}\bar{\theta}/\sqrt{\bar{\alpha}_{t-1}})\,,\quad\hat{\Sigma}_{t}(h)=(\Sigma_{t}^{-1}+\bar{\Sigma}^{-1}/\bar{\alpha}_{t-1})^{-1}\,.$ (8)
@@
The proof is in Section A.2 and comprises four steps. First, we fix $t$ and apply approximation (6). Second, we rewrite the likelihood as in (3). Third, we reparameterize the likelihood, from a function of $s_{t}/\sqrt{\bar{\alpha}_{t}}$ to a function of $s_{t}$. Finally, we combine the likelihood with the Gaussian prior using Lemma 6 in Section A.5. ∎
Algorithm 2 $\tt LaplaceDPS$: Laplace posterior sampling with a diffusion model prior.
@@
4:for stage $t=T,\dots,1$ do $S_{t-1}\sim\mathcal{N}(\hat{\mu}_{t}(S_{t},h),\hat{\Sigma}_{t}(h))$
5:
6:Output: Posterior sample $S_{0}$
The algorithm that samples from the posterior distribution in Theorem 2 is presented in Algorithm 2. We call it Laplace diffusion posterior sampling ($\tt LaplaceDPS$) because its generalization to GLMs utilizes the Laplace approximation. $\tt LaplaceDPS$ works as follows. First, $S_{T}$ is sampled conditioned on evidence $h$ (line 2) from the distribution in (7). This distribution is a product of two Gaussians, one for the prior $\mathcal{N}(\mathbf{0}_{d},I_{d})$ and one for the diffused evidence up to stage $T$, $\mathcal{N}(\sqrt{\bar{\alpha}_{T}}\bar{\theta},\bar{\alpha}_{T}\bar{\Sigma})$. All other variables, $S_{t-1}$ for $t\in[T]$, are sampled conditioned on $s_{t}$ and evidence $h$ (line 3) from the distribution in (8). This distribution is again a product of two Gaussians, one for the conditional prior $\mathcal{N}(\mu_{t}(s_{t}),\Sigma_{t})$ and one for the diffused evidence up to stage $t-1$, $\mathcal{N}(\sqrt{\bar{\alpha}_{t-1}}\bar{\theta},\bar{\alpha}_{t-1}\bar{\Sigma})$. The final output $S_{0}$ is a sample from the diffusion prior.
When compared to Section 2, the prior and empirical evidence are mixed conditionally in a $T$-stage chain. Thus the computational cost of $\tt LaplaceDPS$ is $T$ times higher than that of posterior sampling with a Gaussian prior (Section 2). We validate this empirically in Section 6.2.
4.3 Key Approximation in Theorem 2
Now we motivate our approximation in (6). Using the forward process, $s_{t}=\sqrt{\bar{\alpha}_{t}}s_{0}+\sqrt{1-\bar{\alpha}_{t}}\tilde{\varepsilon}_{t}$ for any $s_{0}$, where $\tilde{\varepsilon}_{t}\sim\mathcal{N}(\mathbf{0}_{d},I_{d})$ is a standard Gaussian noise. This identity can be rearranged as $s_{0}=(s_{t}-\sqrt{1-\bar{\alpha}_{t}}\tilde{\varepsilon}_{t})/\sqrt{\bar{\alpha}_{t}}$. So $s_{0}$ can be viewed as a random variable with mean $s_{t}/\sqrt{\bar{\alpha}_{t}}$, with which we replace it in (6). This approximation is a simplification of Chung et al. [12] where the score component is neglected.
Our approximation has several notable properties. First, $\sqrt{(1-\bar{\alpha}_{t})/\bar{\alpha}_{t}}\to 0$ as $t\to 1$. Therefore, it becomes more precise in later stages of the reverse process. Second, in the absence of evidence $h$, the approximation vanishes, and all posterior distributions in Theorem 2 reduce to the priors in (5). Finally, as the number of observations increases, sampling from the posterior in Theorem 2 is asymptotically consistent.
@@ 4.3 Key Approximation in Theorem 2
Fix $\theta_{*}\in\mathbb{R}^{d}$. Let $\tilde{\theta}\leftarrow{\tt LaplaceDPS}((\mu_{t},\Sigma_{t})_{t\in[T]},h)$, where $h=\left\{(\phi_{\ell},y_{\ell})\right\}_{\ell\in[N]}$ is a history of $N$ observations. Suppose that $\lambda_{d}(\bar{\Sigma}^{-1})\to\infty$ as $N\to\infty$, where $\bar{\Sigma}$ is defined in (3). Then $\mathbb{P}\left(\lim_{N\to\infty}\|\tilde{\theta}-\theta_{*}\|_{2}=0\right)=1$.
@@
The proof is in Section A.3. The key idea is that the conditional posteriors in (7) and (8) concentrate at a scaled unknown model parameter $\theta_{*}$ as the number of observations increases, which we formalize as $\lambda_{d}(\bar{\Sigma}^{-1})\to\infty$. ∎
The bound in Theorem 3 can be interpreted as follows. The sampled parameter $\tilde{\theta}$ approaches the true unknown parameter $\theta_{*}$ as the number of observations $N$ increases. To guarantee that the posterior shrinks uniformly in all directions, we assume that the number of observations in all directions grows linearly with $N$. This is akin to assuming that $\lambda_{d}(\bar{\Sigma}^{-1})=\Omega(N)$. This lower bound can be attained in linear models by choosing feature vectors according to the D-optimal design [38]. Since our proof is asymptotic, we can neglect some finite-time errors and this simplifies the argument.
@@ 4.4 GLM Posterior
The Laplace approximation in GLMs (Section 2.2) naturally generalizes the exact posterior distribution in linear models (Section 2.1). We generalize Theorem 2 to GLMs along the same lines.
@@
Let $p$ be a probability measure over the reverse process (Figure 1). Suppose that we apply approximation (6) for all $t\in[T]$. Then the distribution of $S_{T}\mid h$ is
$p(s_{T}\mid h)\approx\mathcal{N}(s_{T};\hat{\mu}_{T+1}(h),\hat{\Sigma}_{T+1}(h))\,,\quad\hat{\mu}_{T+1}(h)=\sqrt{\bar{\alpha}_{T}}\dot{\theta}_{T+1}\,,\quad\hat{\Sigma}_{T+1}(h)=\bar{\alpha}_{T}\dot{\Sigma}_{T+1}\,,$
where $\dot{\theta}_{T+1},\dot{\Sigma}_{T+1}\leftarrow{\tt IRLS}(\mathbf{0}_{d},I_{d}/\bar{\alpha}_{T},h)$. For $t\in[T]$, the distribution of $S_{t-1}\mid s_{t},h$ is
$p(s_{t-1}\mid s_{t},h)\approx\mathcal{N}(s_{t-1};\hat{\mu}_{t}(s_{t},h),\hat{\Sigma}_{t}(h))\,,\quad\hat{\mu}_{t}(s_{t},h)=\sqrt{\bar{\alpha}_{t-1}}\dot{\theta}_{t}\,,\quad\hat{\Sigma}_{t}(h)=\bar{\alpha}_{t-1}\dot{\Sigma}_{t}\,,$
where $\dot{\theta}_{t},\dot{\Sigma}_{t}\leftarrow{\tt IRLS}(\mu_{t}(s_{t})/\sqrt{\bar{\alpha}_{t-1}},\Sigma_{t}/\bar{\alpha}_{t-1},h)$.
@@
The proof is in Section A.4 and comprises four steps. First, we fix $t$ and apply approximation (6). Second, we reparameterize the prior, from a function of $s_{t}$ to a function of $s_{t}/\sqrt{\bar{\alpha}_{t}}$. Third, we combine the likelihood with the prior using the Laplace approximation. Finally, we repameterize the posterior, from a function of $s_{t}/\sqrt{\bar{\alpha}_{t}}$ to a function of $s_{t}$. ∎
Similarly to Theorem 2, the distributions in Theorem 4 mix evidence with the diffusion model prior. However, this is done implicitly in $\tt IRLS$. The posterior can be sampled from using $\tt LaplaceDPS$, where the mean and covariances would be taken from Theorem 4. Note that Theorem 2 is a special case of Theorem 4 where the mean function $g$ is an identity.
@@ 5 Application to Contextual Bandits
Now we apply our posterior sampling approximations (Section 4) to contextual bandits. A contextual bandit [28, 31] is a classic model for sequential decision making under uncertainty where the agent takes actions conditioned on context. We denote the action set by $\mathcal{A}$ and the context set by $\mathcal{X}$. The mean reward of action $a\in\mathcal{A}$ in context $x\in\mathcal{X}$ is $r(x,a;\theta_{*})$, where $r:\mathcal{X}\times\mathcal{A}\times\Theta\to\mathbb{R}$ is some reward function and $\theta_{*}\in\Theta$ is an unknown model parameter (Section 2). The agent interacts with the bandit for $n$ rounds indexed by $k\in[n]$. In round $k$, the agent observes a context $x_{k}\in\mathcal{X}$, takes an action $a_{k}\in\mathcal{A}$, and observes a stochastic reward $y_{k}=r(x_{k},a_{k};\theta_{*})+\varepsilon_{k}$, where $\varepsilon_{k}$ is independent zero-mean $\sigma^{2}$-sub-Gaussian noise for $\sigma>0$. The goal of the agent is to maximize its cumulative reward in $n$ rounds, or equivalently to minimize its cumulative regret. We define the $n$-round regret as $R(n)=\sum_{k=1}^{n}\mathbb{E}\left[r(x_{k},a_{k,*};\theta_{*})-r(x_{k},a_{k};\theta_{*})\right]$, where $a_{k,*}$ is the optimal action in round $k$.
@@
Arguably the most popular approach to solving contextual bandit problems is Thompson sampling [49, 11, 3]. The key idea in TS is to use the posterior distribution of $\theta_{*}$ to explore. This is done as follows. In round $k$, the model parameter is drawn from the posterior, $\tilde{\theta}_{k}\sim p(\cdot\mid h_{k})$ in (2), where $h_{k}$ is the history of all interactions up to round $k$. After that, the agent takes the action with the highest mean reward under $\tilde{\theta}_{k}$. The pseudo-code of this algorithm is given in Algorithm 3.
@@
A linear bandit [13, 1] is a contextual bandit with a linear reward function $r(x,a;\theta_{*})=\phi(x,a)^{\top}\theta_{*}$, where $\phi:\mathcal{X}\times\mathcal{A}\to\mathbb{R}^{d}$ is a feature extractor. The feature extractor can be non-linear in $x$ and $a$. Therefore, linear bandits can be applied to non-linear functions in $x$ and $a$. The feature extractor can be hand-designed or learned [39]. To simplify notation, we let $\phi_{\ell}=\phi(x_{\ell},a_{\ell})$ be the feature vector of the action in round $k$. So the history of interactions up to round $k$ is $h_{k}=\left\{(\phi_{\ell},y_{\ell})\right\}_{\ell\in[k-1]}$. When the prior distribution is a Gaussian, $p(\theta_{*})=\mathcal{N}(\theta_{*};\theta_{0},\Sigma_{0})$, the posterior in round $k$ is a Gaussian in (3) for $h=h_{k}$. When the prior is a diffusion model, we propose sampling from
$\tilde{\theta}_{k}\leftarrow{\tt LaplaceDPS}((\mu_{t},\Sigma_{t})_{t\in[T]},h_{k})\,,$ (9)
where $\hat{\mu}_{t}$ and $\hat{\Sigma}_{t}$ in $\tt LaplaceDPS$ are set according to Theorem 2. We refer to the resulting algorithm as $\tt DiffTS$.
@@
A generalized linear bandit [16, 23, 32, 25] is an extension of linear bandits to generalized linear models (Section 2.2). When $p(\theta_{*})=\mathcal{N}(\theta_{*};\theta_{0},\Sigma_{0})$, the Laplace approximation to the posterior is a Gaussian (Section 2.2). When the prior distribution is a diffusion model, we propose sampling from (9), where $\hat{\mu}_{t}$ and $\hat{\Sigma}_{t}$ in $\tt LaplaceDPS$ are set according to Theorem 4.

### v2

@@ 4 Posterior Sampling
This section is organized as follows. In Section˜4.1, we show how to sample from a chain of random variables conditioned on observations. In Sections˜4.2 and 4.4, we specialize this to the observation models in Section˜2.
@@ 4.1 Chain Model Posterior
Let $h=\left\{(\phi_{\ell},y_{\ell})\right\}_{\ell\in[N]}$ denote a history of $N$ observations (Section˜2) and $H$ be the corresponding random variable. In this section, we assume that $h$ is fixed. The Markovian structure of the reverse process (Figure˜1) implies that the joint probability distribution conditioned on $h$ factors as
@@
Let $p$ be a probability measure over the reverse process (Figure˜1). Then
@@
The claim is proved in Section˜A.1. ∎
@@ 4.2 Linear Model Posterior
Now we specialize Lemma˜1 to the diffusion model prior (Section˜3) and linear models (Section˜2.1). The prior distribution is the reverse process in (5),
$p(s_{T})=\mathcal{N}(s_{T};\mathbf{0}_{d},I_{d})\,,\quad\forall t\in[T]:p(s_{t-1}\mid s_{t})=\mathcal{N}(s_{t-1};\mu_{t}(s_{t}),\Sigma_{t})\,.$
The term $p(h\mid s_{0})$ is the likelihood of observations in Assumption˜1. The main challenge in using the lemma is that the conditional densities of clean samples $p(s_{0}\mid S_{T})$ and $p(s_{0}\mid s_{t})$ are complex [12]. To get around this, we make an additional assumption, which is discussed in Section˜4.3.
@@
Let $p$ be a probability measure over the reverse process (Figure˜1). Let $\bar{\theta}$ and $\bar{\Sigma}^{-1}$ be defined as in (3). Suppose that
$\textstyle\int_{s_{0}}p(h\mid s_{0})\,p(s_{0}\mid s_{t})\mathop{}\!\mathrm{d}s_{0}\propto p(h\mid s_{t}/\sqrt{\bar{\alpha}_{t}})$ (6)
holds for all $t\in[T]$. Then $p(s_{T}\mid h)\propto\mathcal{N}(s_{T};\hat{\mu}_{T+1}(h),\hat{\Sigma}_{T+1}(h))$, where
$\hat{\mu}_{T+1}(h)=\hat{\Sigma}_{T+1}(h)(\underbrace{I_{d}\,\mathbf{0}_{d}}_{\emph{Prior}}+\underbrace{\bar{\Sigma}^{-1}\bar{\theta}/\sqrt{\bar{\alpha}_{T}}}_{\emph{Evidence}})\,,\quad\hat{\Sigma}_{T+1}(h)=(\underbrace{I_{d}}_{\emph{Prior}}+\underbrace{\bar{\Sigma}^{-1}/\bar{\alpha}_{T}}_{\emph{Evidence}})^{-1}\,.$ (7)
For $t\in[T]$, we have $p(s_{t-1}\mid s_{t},h)\propto\mathcal{N}(s_{t-1};\hat{\mu}_{t}(s_{t},h),\hat{\Sigma}_{t}(h))$, where
$\hat{\mu}_{t}(s_{t},h)=\hat{\Sigma}_{t}(h)(\underbrace{\Sigma_{t}^{-1}\mu_{t}(s_{t})}_{\emph{Prior}}+\underbrace{\bar{\Sigma}^{-1}\bar{\theta}/\sqrt{\bar{\alpha}_{t-1}}}_{\emph{Evidence}})\,,\quad\hat{\Sigma}_{t}(h)=(\underbrace{\Sigma_{t}^{-1}}_{\emph{Prior}}+\underbrace{\bar{\Sigma}^{-1}/\bar{\alpha}_{t-1}}_{\emph{Evidence}})^{-1}\,.$ (8)
@@
The proof is in Section˜A.2. It has four steps. First, we fix stage $t$ and apply approximation (6). Second, we rewrite the likelihood as in (3). Third, we reparameterize it as a function of $s_{t}$. At the end, we combine the likelihood with the Gaussian prior using Lemma˜6 in Section˜A.5. ∎
Algorithm 2 $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt LaplaceDPS$: Laplace posterior sampling with a diffusion model prior.
@@
4:for stage $t=T,\dots,1$ do
5:  $S_{t-1}\sim\mathcal{N}(\hat{\mu}_{t}(S_{t},h),\hat{\Sigma}_{t}(h))$
6:
7:Output: Posterior sample $S_{0}$
The algorithm that samples from the posterior distribution in Theorem˜2 is presented in Algorithm˜2. We call it Laplace diffusion posterior sampling ($\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt LaplaceDPS$) because its generalization to GLMs uses the Laplace approximation. $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt LaplaceDPS$ samples from a chain of products of two distributions: one distribution represents the pre-trained diffusion model and does not depend on history $h$, and the other represents the history $h$. The sampling is implemented as follows. The initial variable $S_{T}$ is sampled conditioned on $h$ (line 2) from the distribution in (7). This distribution is a product of the $h$-independent prior $\mathcal{N}(\mathbf{0}_{d},I_{d})$ and the $h$-dependent distribution of the diffused evidence up to stage $T$, $\mathcal{N}(\sqrt{\bar{\alpha}_{T}}\bar{\theta},\bar{\alpha}_{T}\bar{\Sigma})$. Then, for any $t\in[T]$, $S_{t-1}$ is sampled conditioned on $s_{t}$ and evidence $h$ (line 4) from the distribution in (8). This distribution is a product of the $h$-independent conditional prior $\mathcal{N}(\mu_{t}(s_{t}),\Sigma_{t})$, from the pre-trained model, and the $h$-dependent distribution of the diffused evidence up to stage t1t-1, $\mathcal{N}(\sqrt{\bar{\alpha}_{t-1}}\bar{\theta},\bar{\alpha}_{t-1}\bar{\Sigma})$. The last variable $S_{0}$ is the clean sample. When compared to Section˜2, the prior and evidence are mixed conditionally in a $T$-stage chain. This increases the computational cost $T$ times, as discussed in Section˜8.
4.3 Key Approximation in Theorem˜2
Now we motivate our assumption in (6). Simply put, we assume that $s_{0}=s_{t}/\sqrt{\bar{\alpha}_{t}}$, where $s_{0}$ is a clean sample and $s_{t}$ is the corresponding diffused sample in stage $t$. This is motivated by the forward process, which relates $s_{t}$ and $s_{0}$ as $s_{t}=\sqrt{\bar{\alpha}_{t}}s_{0}+\sqrt{1-\bar{\alpha}_{t}}\varepsilon_{t}$, where $\varepsilon_{t}\sim\mathcal{N}(\mathbf{0}_{d},I_{d})$ is a standard Gaussian noise [19]. After rearranging, we get $s_{0}=(s_{t}-\sqrt{1-\bar{\alpha}_{t}}\varepsilon_{t})/\sqrt{\bar{\alpha}_{t}}$, and therefore $s_{0}$ can be viewed as a random variable with mean $s_{t}/\sqrt{\bar{\alpha}_{t}}$. The consequence of (6) is that the likelihood becomes a function of $s_{t}$, which yields a closed form when multiplied by the conditional prior, also a function of $s_{t}$. Our approximation can be also viewed as the Tweedie’s formula in Chung et al. [12] where the score component is neglected.
Our approximation has several notable properties. First, $\sqrt{(1-\bar{\alpha}_{t})/\bar{\alpha}_{t}}\to 0$ as $t\to 1$. Therefore, it becomes more precise in later stages of the reverse process. Second, in the absence of evidence $h$, the approximation vanishes, and all posterior distributions in Theorem˜2 reduce to the priors in (5). Finally, as the number of observations increases, sampling from the posterior in Theorem˜2 is asymptotically consistent.
@@ 4.3 Key Approximation in Theorem˜2
Fix $\theta_{*}\in\mathbb{R}^{d}$. Let $\tilde{\theta}\leftarrow{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt LaplaceDPS}((\mu_{t},\Sigma_{t})_{t\in[T]},h)$, where $h=\left\{(\phi_{\ell},y_{\ell})\right\}_{\ell\in[N]}$ is a history of $N$ observations. Suppose that $\lambda_{d}(\bar{\Sigma}^{-1})\to\infty$ as $N\to\infty$, where $\bar{\Sigma}$ is defined in (3). Then $\mathbb{P}\left(\lim_{N\to\infty}\|\tilde{\theta}-\theta_{*}\|_{2}=0\right)=1$.
@@
The proof is in Section˜A.3. The key idea is that the conditional posteriors in (7) and (8) concentrate at a scaled unknown model parameter $\theta_{*}$ as the number of observations increases, which we formalize as $\lambda_{d}(\bar{\Sigma}^{-1})\to\infty$. ∎
The bound in Theorem˜3 can be interpreted as follows. The sampled parameter $\tilde{\theta}$ approaches the true unknown parameter $\theta_{*}$ as the number of observations $N$ increases. To guarantee that the posterior shrinks uniformly in all directions, we assume that the number of observations in all directions grows linearly with $N$. This is akin to assuming that $\lambda_{d}(\bar{\Sigma}^{-1})=\Omega(N)$. This lower bound can be attained in linear models by getting observations according to the D-optimal design [39].
@@ 4.4 GLM Posterior
The Laplace approximation in GLMs (Section˜2.2) naturally generalizes the exact posterior distribution in linear models (Section˜2.1). We generalize Theorem˜2 to GLMs along the same lines.
@@
Let $p$ be a probability measure over the reverse process (Figure˜1). Suppose that (6) holds for all $t\in[T]$. Then $p(s_{T}\mid h)\propto\mathcal{N}(s_{T};\hat{\mu}_{T+1}(h),\hat{\Sigma}_{T+1}(h))$, where
$\hat{\mu}_{T+1}(h)=\sqrt{\bar{\alpha}_{T}}\dot{\theta}_{T+1}\,,\quad\hat{\Sigma}_{T+1}(h)=\bar{\alpha}_{T}\dot{\Sigma}_{T+1}\,,\quad\dot{\theta}_{T+1},\dot{\Sigma}_{T+1}\leftarrow{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt IRLS}(\mathbf{0}_{d},I_{d}/\bar{\alpha}_{T},h)\,.$
For $t\in[T]$, we have $p(s_{t-1}\mid s_{t},h)\propto\mathcal{N}(s_{t-1};\hat{\mu}_{t}(s_{t},h),\hat{\Sigma}_{t}(h))$, where
$\hat{\mu}_{t}(s_{t},h)=\sqrt{\bar{\alpha}_{t-1}}\dot{\theta}_{t}\,,\quad\hat{\Sigma}_{t}(h)=\bar{\alpha}_{t-1}\dot{\Sigma}_{t}\,,\quad\dot{\theta}_{t},\dot{\Sigma}_{t}\leftarrow{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt IRLS}(\mu_{t}(s_{t})/\sqrt{\bar{\alpha}_{t-1}},\Sigma_{t}/\bar{\alpha}_{t-1},h)\,.$
@@
The proof is in Section˜A.4. It has four steps. First, we fix stage $t$ and apply approximation (6). Second, we reparameterize the prior, from a function of $s_{t}$ to a function of $s_{t}/\sqrt{\bar{\alpha}_{t}}$. Third, we combine the likelihood with the prior using the Laplace approximation. Finally, we repameterize the posterior, from a function of $s_{t}/\sqrt{\bar{\alpha}_{t}}$ to a function of $s_{t}$. ∎
Similarly to Theorem˜2, the distributions in Theorem˜4 mix evidence with the diffusion model prior. However, this is done implicitly in $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt IRLS$. The posterior can be sampled from using $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt LaplaceDPS$, where the mean and covariances would be taken from Theorem˜4. Note that Theorem˜2 is a special case of Theorem˜4 where the mean function $g$ is an identity.
@@ 5 Application to Contextual Bandits
Now we apply our posterior sampling approximations (Section˜4) to contextual bandits. A contextual bandit [29, 32] is a classic model for sequential decision making under uncertainty where the agent takes actions conditioned on context. We denote the action set by $\mathcal{A}$ and the context set by $\mathcal{X}$. The mean reward for taking action $a\in\mathcal{A}$ in context $x\in\mathcal{X}$ is $r(x,a;\theta_{*})$, where $r:\mathcal{X}\times\mathcal{A}\times\Theta\to\mathbb{R}$ is a reward function and $\theta_{*}\in\Theta$ is a model parameter (Section˜2). The agent interacts with the bandit for $n$ rounds indexed by $k\in[n]$. In round $k$, it observes a context $x_{k}\in\mathcal{X}$, takes an action $a_{k}\in\mathcal{A}$, and observes its stochastic reward $y_{k}=r(x_{k},a_{k};\theta_{*})+\varepsilon_{k}$ with independent noise $\varepsilon_{k}$. We assume that the noise is zero-mean $\sigma^{2}$-sub-Gaussian for $\sigma>0$. The objective of the agent is to maximize its cumulative reward in $n$ rounds, or equivalently to minimize its cumulative regret. We define the $n$-round regret as
$\textstyle R(n)=\sum_{k=1}^{n}\mathbb{E}\left[r(x_{k},a_{k,*};\theta_{*})-r(x_{k},a_{k};\theta_{*})\right]\,$, (9)
where $a_{k,*}=\operatorname*{arg\,max\,}_{a\in\mathcal{A}}r(x_{k},a;\theta_{*})$ is the optimal action in round $k$.
@@
Arguably the most popular method for solving contextual bandit problems is Thompson sampling [50, 11, 3]. The key idea in TS is to use the posterior distribution of $\theta_{*}$ to explore. This is done as follows. In round $k$, the model parameter is drawn from the posterior in (2), $\tilde{\theta}_{k}\sim p(\cdot\mid h_{k})$, where $h_{k}$ is the history of all interactions up to round $k$. After that, the agent takes the action with the highest mean reward under $\tilde{\theta}_{k}$. The pseudo-code of this algorithm is given in Algorithm˜3.
@@
A linear bandit [13] has a linear reward function $r(x,a;\theta_{*})=\phi(x,a)^{\top}\theta_{*}$, where $\phi:\mathcal{X}\times\mathcal{A}\to\mathbb{R}^{d}$ is a feature extractor. The feature extractor can be non-linear in $x$ and $a$. Therefore, linear bandits can be applied to non-linear functions of $x$ and $a$. The feature extractor can be either learned [40] or hand-crafted. We denote the feature vector of the action in round $k$ by $\phi_{k}=\phi(x_{k},a_{k})$. Therefore, the history of interactions up to round $k$ is $h_{k}=\left\{(\phi_{\ell},y_{\ell})\right\}_{\ell\in[k-1]}$. When the prior distribution is a Gaussian, $p(\theta_{*})=\mathcal{N}(\theta_{*};\theta_{0},\Sigma_{0})$, the posterior in round $k$ is a Gaussian in (3) for $h=h_{k}$. When the prior is a diffusion model, we propose sampling from the posterior using
$\tilde{\theta}_{k}\leftarrow{\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt LaplaceDPS}((\mu_{t},\Sigma_{t})_{t\in[T]},h_{k})\,$, (10)
where $\hat{\mu}_{t}$ and $\hat{\Sigma}_{t}$ in $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt LaplaceDPS$ are computed as in Theorem˜2. We call this algorithm $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt DiffTS$.
@@
A generalized linear bandit [17, 24, 33, 26] is an extension of linear bandits to generalized linear models (Section˜2.2). When $p(\theta_{*})=\mathcal{N}(\theta_{*};\theta_{0},\Sigma_{0})$, the Laplace approximation to the posterior is a Gaussian (Section˜2.2). When the prior is a diffusion model, we propose posterior sampling using (10), where $\hat{\mu}_{t}$ and $\hat{\Sigma}_{t}$ in $\color[rgb]{0,1,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,1,0}\pgfsys@color@cmyk@stroke{1}{0}{1}{0}\pgfsys@color@cmyk@fill{1}{0}{1}{0}\tt LaplaceDPS$ are computed as in Theorem˜4.